6) The sound from the whistle of a truck moving at 17.0 m/s toward a cliff reflects from the cliff and is heard by the driver of the truck. When it is parked at the side of the road, the whistle from the truck produces sound of frequency 2.267 kHz. As observed by the driver of the truck, what is the change in frequency the sound from the whistle due to the reflection? The speed of sound is 343 m/s. A) 236 Hz B) 228 Hz C) 207 Hz
Answer #1

Frequency heard of reflected sound f2 = f0*v/(v-vs) *(v+vs)/v

= 2267*(343+17)/(343-17) = 2503 Hz

change = 2503 - 2267 = 236 Hz option A answer